OpenStudy (anonymous):

How do I solve this??: The scale on a map is approx. 5/8 inches to 15 miles. If the distance from city A to city B is 9.5 inches on the map, how far apart are the two cities?

OpenStudy (ash2326):

It's given that 5/8 inches in map are equal to 15 miles so 1 inch equals= 15*8/5 miles= 24 miles Distance between A and B is 9.5 inches on the map so they are 9.5*24 miles apart= 228 miles

OpenStudy (anonymous):

thank you so much!!

OpenStudy (ash2326):

you're welcome Laurenn. Did you understand?

OpenStudy (paxpolaris):

\[{5/8\ inches \over 15\ miles}={1inch \over 24 miles}\] \[\Large {1\ \cancel{inch}\over 24\ miles}={9.5\ \cancel{inches}\over x}\]

OpenStudy (anonymous):

Yes, I understand! Thank you very much!!